# Robust purely optical signatures of Floquet-Bloch states in crystals

### v1

@@ Acknowledgements.
- Lucchini et al. [2016] M. Lucchini, S. A. Sato, A. Ludwig, J. Herrmann, M. Volkov, L. Kasmi, Y. Shinohara, K. Yabana, L. Gallmann, and U. Keller, Attosecond dynamical franz-keldysh effect in polycrystalline diamond, Science 353, 916 (2016).
- Zhang et al. [2021] Y. Zhang, T. Huang, L. Li, P. Lan, and P. Lu, Intensity and wavelength dependence of anisotropic nonlinear absorption inside mgo, Opt. Quantum Electron 53, 158 (2021).
- Reislöhner et al. [2023] J. Reislöhner, X. Chen, D. Kim, S. Botti, and A. N. Pfeiffer, Dynamical franz-keldysh effect in diamond in the deep ultraviolet probed by transient absorption and dispersion spectroscopy using a miniature beamline, Phys. Rev. Lett. 131, 136902 (2023).
@@
- Otobe et al. [2016] T. Otobe, Y. Shinohara, S. A. Sato, and K. Yabana, Femtosecond time-resolved dynamical franz-keldysh effect, Phys. Rev. B 93, 045124 (2016).
- Sato et al. [2018] S. A. Sato, H. Hübener, U. De Giovannini, and A. Rubio, Ab initio simulation of attosecond transient absorption spectroscopy in two-dimensional materials, Appl. Sci. 8, 1777 (2018).
- Du et al. [2019] M. Du, C. Liu, Y. Zheng, Z. Zeng, and R. Li, Attosecond transient-absorption spectroscopy in one-dimensional periodic crystals, Phys. Rev. A 100, 043840 (2019).
- Tancogne-Dejean et al. [2020] N. Tancogne-Dejean, M. A. Sentef, and A. Rubio, Ultrafast transient absorption spectroscopy of the charge-transfer insulator NiO: Beyond the dynamical franz-keldysh effect, Phys. Rev. B 102, 115106 (2020).
- Dong and Liu [2022] F. Dong and J. Liu, Fishbone resonance structure in the attosecond transient absorption spectrum of graphene, Phys. Rev. A 106, 063107 (2022).
- Mizumoto and Kayanuma [2005] Y. Mizumoto and Y. Kayanuma, Band renormalization of semiconductors by high-intensity infrared laser: One-dimensional model, Phys. Rev. B 72, 115203 (2005).
- Mizumoto et al. [2006] Y. Mizumoto, Y. Kayanuma, A. Srivastava, J. Kono, and A. H. Chin, Dressed-band theory for semiconductors in a high-intensity infrared laser field, Phys. Rev. B 74, 045216 (2006).
- Kim [2023] Y. Kim, Ultrafast transient absorption spectroscopy of strongly correlated mott insulators, Phys. Rev. B 108, 035122 (2023).
- Tiwari and Franco [2025] V. Tiwari and I. Franco, First-principle based floquet engineering of solids in the velocity gauge (2025), https://arxiv.org/abs/2501.15612.
@@
- Passos et al. [2018] D. J. Passos, G. B. Ventura, J. M. V. P. Lopes, J. M. B. L. dos Santos, and N. M. R. Peres, Nonlinear optical responses of crystalline systems: Results from a velocity gauge analysis, Phys. Rev. B 97, 235446 (2018).
@@
- [52] See attached Supplemental Material for ZnO electronic structure computation, two-band model parameters and optical response of laser-dressed two-band model.
@@
- Bagnall et al. [1997] D. M. Bagnall, Y. F. Chen, Z. Zhu, T. Yao, S. Koyama, M. Y. Shen, and T. Goto, Optically pumped lasing of zno at room temperature, Appl. Phys. Lett. 70, 2230 (1997).
@@
- Kruchinin et al. [2018] S. Y. Kruchinin, F. Krausz, and V. S. Yakovlev, Colloquium : Strong-field phenomena in periodic systems, Rev. Mod. Phys. 90, 021002 (2018).

### v2

@@ Acknowledgements.
- Kubo [1963] R. Kubo, Stochastic liouville equations, Journal of Mathematical Physics 4, 174–183 (1963).
- Kubo [1957] R. Kubo, Statistical-mechanical theory of irreversible processes. i. general theory and simple applications to magnetic and conduction problems, Journal of the Physical Society of Japan 12, 570–586 (1957).
- Mukamel [1995] S. Mukamel, Principles of Nonlinear Optical Spectroscopy, Oxford series in optical and imaging sciences (Oxford University Press, 1995).
- Tiwari and Franco [2025] V. Tiwari and I. Franco, First-principle based floquet engineering of solids in the velocity gauge, To be published (2025).
@@
- Ghimire et al. [2010] S. Ghimire, A. D. DiChiara, E. Sistrunk, P. Agostini, L. F. DiMauro, and D. A. Reis, Observation of high-order harmonic generation in a bulk crystal, Nature Physics 7, 138–141 (2010).
@@
- [44] See Supplemental Material at [url] for ZnO electronic structure computation, two-band model parameters, optical response of laser-dressed two-band model, optical response of laser-dressed Hubbard-Holstein model, and absorption spectra of laser-dressed ZnO with varying spectral broadening.
- Szidarovszky et al. [2019] T. Szidarovszky, A. G. Csaszar, G. J. Halasz, and A. Vibok, Rovibronic spectra of molecules dressed by light fields, Physical Review A 100, 10.1103/physreva.100.033414 (2019).
@@
- Ozgur et al. [2005] U. Ozgur, Y. I. Alivov, C. Liu, A. Teke, M. A. Reshchikov, S. Dogan, V. Avrutin, S.-J. Cho, and H. Morkoc, A comprehensive review of zno materials and devices, Journal of Applied Physics 98, 10.1063/1.1992666 (2005).
- Rohlfing and Louie [2000] M. Rohlfing and S. G. Louie, Electron-hole excitations and optical spectra from first principles, Physical Review B 62, 4927–4944 (2000).
@@
- Autler and Townes [1955] S. H. Autler and C. H. Townes, Stark effect in rapidly varying fields, Physical Review 100, 703–722 (1955).
- Kim et al. [2020] J. Kim, J. S. Lim, H.-R. Noh, and S. K. Kim, Experimental observation of the autler–townes splitting in polyatomic molecules, The Journal of Physical Chemistry Letters 11, 6791–6795 (2020).
- Schüler et al. [2018] M. Schüler, M. Eckstein, and P. Werner, Truncating the memory time in nonequilibrium dynamical mean field theory calculations, Physical Review B 97, 10.1103/physrevb.97.245129 (2018).
- Johnston et al. [2010] S. Johnston, F. Vernay, B. Moritz, Z.-X. Shen, N. Nagaosa, J. Zaanen, and T. P. Devereaux, Systematic study of electron-phonon coupling to oxygen modes across the cuprates, Physical Review B 82, 10.1103/physrevb.82.064513 (2010).
- Esslinger [2010] T. Esslinger, Fermi-hubbard physics with atoms in an optical lattice, Annual Review of Condensed Matter Physics 1, 129–152 (2010).
- Fehske et al. [2004] H. Fehske, G. Wellein, G. Hager, A. Weiße, and A. R. Bishop, Quantum lattice dynamical effects on single-particle excitations in one-dimensional mott and peierls insulators, Physical Review B 69, 10.1103/physrevb.69.165115 (2004).
- Hu et al. [2018] W. Hu, B. Gu, and I. Franco, Lessons on electronic decoherence in molecules from exact modeling, The Journal of Chemical Physics 148, 10.1063/1.5004578 (2018).
- Light and Carrington [2007] J. C. Light and T. Carrington, Discrete-variable representations and their utilization, in Advances in Chemical Physics (John Wiley & Sons, Inc., 2007) pp. 263–310.
- Jordan and Wigner [1928] P. Jordan and E. Wigner, Über das Paulische Äquivalenzverbot, Z. Phys. 47, 631–651 (1928).
- Colbert and Miller [1992] D. T. Colbert and W. H. Miller, A novel discrete variable representation for quantum mechanical reactive scattering via the s-matrix kohn method, The Journal of Chemical Physics 96, 1982–1991 (1992).